OpenStudy (anonymous):

Solve the equation. 2p=2 p= q-3=7 q= Help really needed.

undeadknight26 (undeadknight26):

2p = 2 /2 /2 p =1. try the second one by your self :D

OpenStudy (anonymous):

2p=2 divide both sides by 2 p = 1 q-3=8 add 3 to both sides q = 11

undeadknight26 (undeadknight26):

Oh yeah what he said ↑↑↑

OpenStudy (anonymous):

p=1 q=8

undeadknight26 (undeadknight26):

no guys q = 10: q-3=7 +3 +3 q = 10

OpenStudy (anonymous):

Yea what he said. It's 10 - don't know why i put 11.

undeadknight26 (undeadknight26):

you made a typo and put q - 3 = "8" not 7 XP

OpenStudy (anonymous):

Oh I see, I accidently typed on my equation. I put 8 instead of 7.

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

So p is 1 and q is 10 ?

undeadknight26 (undeadknight26):

precisely :D

OpenStudy (anonymous):

Ok thanks XD